# Loop Feynman integration on a quantum computer

### v1

@@ Abstract
This Letter investigates in detail the performance and advantages of a new quantum Monte Carlo integrator, dubbed Quantum Fourier Iterative Amplitude Estimation (QFIAE), to numerically evaluate for the first time loop Feynman integrals in a near-term quantum computer and a quantum simulator. In order to achieve a quadratic speedup, QFIAE introduces a Quantum Neural Network (QNN) that efficiently decomposes the multidimensional integrand into its Fourier series. For a one-loop tadpole Feynman diagram, we have successfully implemented the quantum algorithm on a real quantum computer and obtained a reasonable agreement with the analytical values. One-loop Feynman diagrams with more external legs have been analyzed in a quantum simulator. These results thoroughly illustrate how our quantum algorithm effectively estimates loop Feynman integrals and the method employed could also find applications in other fields such as finance, artificial intelligence, or other physical sciences.
@@ INTRODUCTION
To unravel the mysteries of the universe at its most fundamental level, Quantum Field Theory (QFT) has proven to be an astounding theory, showcasing significant agreement between theoretical predictions and experimental observations at particle colliders. However, as our understanding of the particle realm advances, the necessity to explore higher and higher energies poses significant challenges for precise measurements and precise theoretical predictions. In this regard, the perturbative approach in QFT has gained immense importance, emerging as a crucial framework for deriving accurate predictions in high-energy physics. Nevertheless, computing higher-order contributions in perturbative QFT is far from straightforward. The primary challenge lies in dealing with virtual quantum fluctuations, leading to the intricacies of multiloop-multileg Feynman integrals. The difficulty in computing these integrals arises from their multidimensional nature, dependence on multiple scales, and the presence of ultraviolet (UV), infrared (IR), and threshold singularities. Tackling these challenges often involves tiresome regularization, renormalization, and subtraction techniques
@@
Considering the considerable computational demands of standard methods, the imperative to seek alternative approaches becomes evident. This necessity arises from the need to overcome existing state-of-the-art limitations. In response, there is a growing interest in exploring innovative strategies rooted in quantum computing to address traditionally challenging problems spanning diverse domains. The potential acceleration offered by quantum computers has spurred various ideas, including Grover’s algorithm for efficient database querying Grover (1997), Shor’s algorithm for factorization of large integers Shor (1997), and quantum annealing for Hamiltonian minimization Kadowaki and Nishimori (1998). In the realm of particle physics, quantum algorithms have made inroads into different areas, such as lattice gauge theories Zohar et al. (2015); Jordan et al. (2012). Moreover, these algorithms have been deployed in various tasks related to high-energy colliders Delgado et al. (2022). These applications encompass jet identification and clustering Wei et al. (2020); Delgado and Thaler (2022); Martínez de Lejarza et al. (2022), determination of parton densities (PDFs) Pérez-Salinas et al. (2021), simulation of parton showers Bepari et al. (2022), anomaly detection Woźniak et al. (2023); Schuhmacher et al. (2023); Bermot et al. (2023), integration of elementary particle processes Agliardi et al. (2022), and the establishment of the causal structure of multiloop Feynman diagrams Ramírez-Uribe et al. (2022); Clemente et al. (2023). The rapid expansion of applications underscores the versatility of quantum algorithms for a wide array of purposes Meglio et al. (2023).
@@
In light of these recent achievements, this Letter seeks to investigate the potential of quantum algorithms for the efficient computation of loop Feynman integrals. Specifically, we delve into the capabilities of Quantum Fourier Iterative Amplitude Estimation (QFIAE), a quantum algorithm recently introduced in de Lejarza et al. (2023); de Lejarza (2023). QFIAE serves as an end-to-end quantum Monte Carlo integrator, showcasing the targeted potential for a quadratic speedup.
@@
The workflow of QFIAE, which is illustrated in Fig. 1, initiates by decomposing the target function into its Fourier series using a Quantum Neural Network (QNN) following a data re-uploading approach Schuld et al. (2021); Casas and Cervera-Lierta (2023); Pérez-Salinas et al. (2020). Subsequently, each trigonometric term of the Fourier series undergoes quantum integration using Iterative Quantum Amplitude Estimation (IQAE) Grinko et al. (2021), an efficient version of Quantum Amplitude Estimation(QAE) Brassard et al. (2002).
@@
The Fourier decomposition enables the target function to be encodable with a minimum number of quantum arithmetic operations. At the same time, it takes advantage of the fact that the sine function is more suitable for integration in a quantum approach. The QNN is the central ingredient of QFIAE that offers a viable strategy to retain the quadratic speedup with respect to other quantum integration algorithms recently proposed, such as Fourier Quantum Monte Carlo Integration (FQMCI) Herbert (2022). The key idea of FQMCI is also to use Fourier series decomposition to approximate the integrand and then estimate each component separately using QAE. Nevertheless, it relies on certain assumptions regarding the acquisition of the Fourier coefficients, which may not hold in general. When these assumptions are not met, the quantum speedup might be wiped out. The QNN ensures a reliable extraction of the Fourier coefficients in a quantum way.
@@
The second crucial aspect of QFIAE involves leveraging the advantages of IQAE over QAE. Quantum Amplitude Estimation (QAE) Brassard et al. (2002) is a quantum algorithm that estimates the amplitude value of a quantum state using amplitude amplification, a generalization of Grover’s querying algorithm Grover (1997). This process enhances the probability of measuring a desired state over a nondesired state. However, QAE has inherent limitations, including its reliance on the resource-intensive Quantum Phase Estimation (QPE) subroutine Nielsen and Chuang (2011), which involves operations considered computationally expensive for current Noisy Intermediate Scale Quantum (NISQ) devices. In certain scenarios, these limitations may compromise the anticipated quadratic speedup promised by QAE. To address this challenge, IQAE replaces the QPE subroutine with a classically efficient post-processing method, reducing the required number of qubits and gates while maintaining the asymptotic quadratic speedup.
@@
I QUANTUM INTEGRATION OF LOOP FEYNMAN INTEGRALS
@@
Among other advantages, this transformation provides a more intuitive understanding of the singular structure of loop integrals Buchta et al. (2014); Aguilera-Verdugo et al. (2019). In particular, the most remarkable property of LTD is the existence of a manifestly causal representation ”Aguilera-Verdugo et al. (2020); Aguilera-Verdugo et al. (2021); Ramírez-Uribe et al. (2021); Jesús Aguilera-Verdugo et al. (2021); Sborlini (2021); Torres Bobadilla (2021), i.e., an integrand representation where certain nonphysical singularities are absent and therefore yields integrands that are numerically more stable. The fact that the integration domain in LTD is Euclidean and not Minkowski also brings additional advantages, both for analytic applications such as asymptotic expansions Driencourt-Mangin et al. (2018); Plenter and Rodrigo (2021), where the hierarchy of scales is well defined, and numerical applications Buchta et al. (2017); Driencourt-Mangin et al. (2022) because the number of loop integration variables is independent of the number of external particles. For example, at one loop the number of independent integration variables is always three, the number of spatial components of the loop momentum, although for certain kinematic configurations this number can be reduced when the dependence on any of these variables is trivial. This is the case for tadpole and bubble diagrams at one loop. Moreover, LTD offers a unified framework for cross-section calculations, since the dual representation of loop integrals in Euclidean domains allows a direct combination of virtual and real contributions at the integrand level, resulting in a fully local cancellation of IR Hernandez-Pinto et al. (2016); Sborlini et al. (2016a, b); Prisco and Tramontano (2021); de Jesús Aguilera-Verdugo et al. (2021) and UV Driencourt-Mangin et al. (2018, 2019) singularities without the need for modifying the dimensions of space-time, such as in Dimensional Regularization (DREG) Bollini and Giambiagi (1972); ’t Hooft and Veltman (1972).
@@
Tadpole loop integral: the first Feynman loop integral we address is the one-loop tadpole integral
@@
is well defined in the four physical dimensions of space-time, where the integration measure reads $\int_{\boldsymbol{\ell}}=\int d^{3}\boldsymbol{\ell}/(2\pi)^{3}$. In Eq. (2), we have defined the on-shell energies $q_{1,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m^{2}-\imath 0}$ and $q_{{\rm UV},0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+\mu_{\rm UV}^{2}-\imath 0}$, where $\mu_{\rm UV}$ is the renormalization scale. The tadpole integral is in fact a one-dimensional integral because the integrand is independent of the solid angle, and only depends on the modulus of the loop three-momentum.
In order to implement QFIAE in a real quantum computer we must consider that quantum computers operating in the NISQ era encounter diverse sources of noise, ranging from quantum effects like decoherence to hardware-specific errors, including gate, readout, and calibration errors. At the same time, working on the same hardware technology (superconducting qubits), we advocate for a hardware-agnostic implementation strategy. Specifically, we propose an approach where the two modules of our quantum algorithm are implemented on two different quantum computers supplied by different providers. This enables a minimization of the impact of hardware-specific noise on the overall algorithmic performance.
@@
On the other hand, a 5-qubit IQAE algorithm has been executed using Qiskit Qiskit contributors (2023) on the 27-qubit IBMQ superconducting device ibmq_mumbai. To mitigate quantum noise during execution, we employed a pulse-efficient transpilation technique Earnest et al. (2021). Effectively reducing the number of two-qubit gates by harnessing the hardware-native cross-resonance interaction. While it has previously demonstrated promise in the context of Variational Quantum Circuits (VQC) Melo et al. (2023), it has not yet been extensively explored for fault-tolerant applications. Considering the positive results presented in Table S2, this work serves as an intriguing starting point for the applicability of this mitigation technique to fault-tolerant algorithms. Furthermore, we also applied two more error mitigation techniques, Dynamical Decoupling (DD) and Zero Noise Extrapolation (ZNE), which are automatized within the Qiskit Runtime Estimator primitive IBMQ (2023).
Fourier IQAE
${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$
$m=5,\mu_{\rm UV}=m/2$
${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$
$m=5,\mu_{\rm UV}=2m$
@@
Table 1: Renormalized tadpole integral ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$ on Qibo and IBMQ, as a function of the ratio of the renormalization scale $\mu_{\rm UV}$ to the mass $m$, which is fixed to $m=5$. The Fourier decomposition and IQAE integration are performed either with a quantum simulator (S) or on a real quantum device (Q).
In particular, the hardware implementation has been done by fixing the mass to $m=5$ GeV, and considering two values of the renormalization scale, $\mu_{\rm UV}=2m$ and $\mu_{\rm UV}=m/2$. The uncertainties on the integrals presented in Table S2 have been computed as the quadratic sum of the individual uncertainties provided by IQAE for each trigonometric component. These uncertainties are statistical in nature and are to be combined with the statistical uncertainties from the noise of the quantum devices employed in both the QNN training and IQAE integration modules of the QFIAE algorithm. Furthermore, there are also systematic uncertainties from the classical or quantum Fourier series truncation of the integrand function ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$. Taking into account all statistical and systematic uncertainties the obtained results are in agreement with the analytical values within uncertainties.
The results presented in Table S2 show a relatively small deviation from the analytical value for both $\mu_{\rm UV}=m/2$ and $\mu_{\rm UV}=2m$. In particular, the deviations when both parts of the algorithm are executed on a quantum computer (last row of Table S2) are $4.4\%$ and $5.9\%$ for $\mu_{\rm UV}=m/2$ and $\mu_{\rm UV}=2m$, respectively. Although this relative error might seem large at first glance, it represents a noteworthy achievement for the current state of quantum computing technology. Moreover, it is the first application of an end-to-end quantum algorithm executed on a quantum computer for estimating loop Feynman integrals.
Bubble loop integral: QFIAE is also effective in dealing with Feynman loop integrals with threshold singularities. Therefore, we consider the one-loop bubble integral:
@@
${\cal A}^{(1)}_{2}(p,m_{1},m_{2})=\int_{\boldsymbol{\ell}}\left[\frac{1}{x_{2}}\left(\frac{1}{\lambda^{+}}+\frac{1}{\lambda^{-}}\right)-\frac{1}{4(q_{{\rm UV},0}^{(+)})^{3}}\right]\leavevmode\nobreak\ ,$ (4)
where $x_{2}=\prod_{i=1,2}2q_{i,0}^{(+)}$, $\lambda^{\pm}=\sum_{i=1,2}q_{i,0}^{(+)}\pm p_{0}$, and the on-shell energies are given by $q_{i,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m_{i}^{2}-\imath 0}$ with $i\in\{1,2\}$, assuming the external momentum has vanishing spatial components, $p=(p_{0},{\bf 0})$. If $p_{0}^{2}<(m_{1}+m_{2})^{2}$ the integral is purely real. Otherwise, it gets an imaginary contribution from the unitary threshold singularity at $\lambda^{-}\to 0$, assuming $p_{0}>0$. To deal with this threshold singularity, it is convenient to introduce a contour deformation in the complex plane, which will smooth the behavior of the function in the vicinity of the threshold without altering the result of the integral. We consider $m_{1}=m_{2}=m$, and proceed to estimate the integral using two different QNNs to fit separately the real and imaginary parts of the integrand and then integrate each Fourier series using IQAE. We have implemented QFIAE on two different quantum simulators. For the QNNs, we employed Pennylane Bergholm et al. (2018) whereas for IQAE, we utilized Qibo.
@@
Triangle loop integral: we now consider the one-loop three-point function, which corresponds to a triangle loop topology. The LTD representation of this integral is given by Aguilera-Verdugo et al. (2021)
@@
with $x_{3}=\prod_{i=1}^{3}2q_{i,0}^{(+)}$, where now the on-shell energies are $q_{1,0}^{(+)}=\sqrt{(\boldsymbol{\ell}+{\bf p}_{1})^{2}+m_{1}^{2}-\imath 0}$, and $q_{i,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m_{i}^{2}-\imath 0}$ for $i\in\{2,3\}$. We work in the centre of mass (cms) frame where ${\bf p}_{12}={\bf p}_{1}+{\bf p}_{2}=0$, $p_{12,0}=p_{1,0}+p_{2,0}=\sqrt{s}$ and the external momenta $p_{1}$ and $p_{2}$ are back-to-back along the $z$ axis. The causal denominators are
@@ II CONCLUSIONS
In this Letter, we have for the first time successfully implemented a quantum algorithm for evaluating Feynman loop integrals in quantum hardware. Our approach is based on the nonconventional LTD representation of Feynman loop integrals and the QFIAE quantum algorithm which provides a theoretical quadratic quantum speedup.
We have provided a clear indication of how the optimization of a quantum circuit, at the level of pulses and the application of combined error mitigation techniques, enable the collection of accurate results.
@@
Specifically, we have employed QFIAE to numerically integrate the one-dimensional tadpole Feynman loop integral entirely on real quantum hardware provided by Qibo and IBM. Moreover, loop Feynman integrals with more external legs and dimensions have also been integrated using QFIAE on quantum simulators across various kinematical regions, circumventing different threshold singularities using a contour deformation. The obtained results demonstrate the remarkable performance of this quantum algorithm in estimating these integrals.
@@
Overall, our study highlights the potential of Quantum Monte Carlo integrators, particularly the QFIAE quantum algorithm, in integrating loop Feynman integrals. These findings, some of them obtained from quantum superconducting devices, serve as an initial step toward the search for a quantum method capable of speeding up the computationally expensive task of evaluating multiloop-multileg Feynman integrals. Moreover, they offer a proof of concept for the method’s applicability to real-world challenges across various domains that require Monte Carlo integration, such as finance, artificial intelligence or other physical sciences.
@@
Still, a technical challenge lies in identifying a suitable Ansatz capable of ensuring the trainability of multidimensional functions, which is essential for addressing higher-loop integrals. While universality has proven effective in lower dimensions, extending it beyond one dimension remains an open problem. Although our study achieved a successful fit for a two-dimensional function, the search for a viable Ansatz in higher dimensions poses significant technical challenges, since it involves a tradeoff between trainability and expressivity of the variational quantum circuit. Progressing in the previous limitation would assure, according to the algorithm proposed in this Letter, a significant noteworthy advancement over classical Monte Carlo methods for computing these integrals, presenting new opportunities for unlocking previously inaccessible levels of precision in theoretical predictions for high-energy physics at colliders.

### v2

@@ Abstract
This work investigates in detail the performance and advantages of a new quantum Monte Carlo integrator, dubbed Quantum Fourier Iterative Amplitude Estimation (QFIAE), to numerically evaluate for the first time loop Feynman integrals in a near-term quantum computer and a quantum simulator. In order to achieve a quadratic speedup, QFIAE introduces a Quantum Neural Network (QNN) that efficiently decomposes the multidimensional integrand into its Fourier series. For a one-loop tadpole Feynman diagram, we have successfully implemented the quantum algorithm on a real quantum computer and obtained a reasonable agreement with the analytical values. One-loop Feynman diagrams with more external legs have been analyzed in a quantum simulator. These results thoroughly illustrate how our quantum algorithm effectively estimates loop Feynman integrals and the method employed could also find applications in other fields such as finance, artificial intelligence, or other physical sciences.
@@
To unravel the mysteries of the universe at its most fundamental level, Quantum Field Theory (QFT) stands out as an astounding theory, demonstrating significant agreement between theoretical predictions and experimental observations at particle colliders. However, as our comprehension of the particle realm progresses, the need to explore higher energies presents substantial challenges for precise measurements and theoretical predictions. In this regard, the perturbative approach in QFT has gained immense importance, emerging as a crucial framework for deriving accurate predictions in high-energy physics. Nevertheless, computing higher-order contributions in perturbative QFT is far from straightforward. The primary challenge lies in dealing with virtual quantum fluctuations, leading to the intricacies of multiloop-multileg Feynman integrals. The difficulty stems from their multidimensional nature, dependence on multiple scales, and the presence of ultraviolet (UV), infrared (IR), and threshold singularities. Tackling these challenges often involves tiresome regularization, renormalization, and subtraction techniques.
@@
Considering the enormous computational demands of standard methods, the imperative to seek alternative approaches becomes evident to overcome existing state-of-the-art limitations. In response, there is a growing interest in exploring innovative strategies rooted in quantum computing to address traditionally challenging problems spanning diverse domains. The potential acceleration offered by quantum computers has spurred various ideas, including Grover’s algorithm for efficient database querying Grover (1997), Shor’s algorithm for factorization of large integers Shor (1997), and quantum annealing for Hamiltonian minimization Kadowaki and Nishimori (1998). In the realm of particle physics, quantum algorithms have made inroads into different areas, such as lattice gauge theories Zohar et al. (2015); Jordan et al. (2012). Moreover, these algorithms have been deployed in various tasks related to high-energy colliders Delgado et al. (2022). These applications encompass jet identification and clustering Wei et al. (2020); Delgado and Thaler (2022); Martínez de Lejarza et al. (2022); de Lejarza et al. (2022), determination Pérez-Salinas et al. (2021) and integration Cruz-Martinez et al. (2023) of parton densities, simulation of parton showers Bepari et al. (2022), anomaly detection Woźniak et al. (2023); Schuhmacher et al. (2023); Bermot et al. (2023), integration of elementary particle processes Agliardi et al. (2022), calculation of color factors in QCD Chawdhry and Pellen (2023), and the establishment of the causal structure of multiloop Feynman diagrams Ramírez-Uribe et al. (2022); Clemente et al. (2023). The rapid expansion of applications underscores the versatility of quantum algorithms for a wide array of purposes Meglio et al. (2023).
@@
In light of these recent achievements, this article seeks to investigate the potential of quantum algorithms for the efficient computation of loop Feynman integrals. Specifically, we delve into the capabilities of Quantum Fourier Iterative Amplitude Estimation (QFIAE), a quantum algorithm recently introduced in de Lejarza et al. (2023); de Lejarza (2023). QFIAE serves as an end-to-end quantum Monte Carlo integrator, showcasing the aimed potential for a quadratic speedup in querying the probability distribution function that follows the target function.
@@
The workflow of QFIAE, illustrated in Fig. 1, initiates by decomposing the target function into its Fourier series using a Quantum Neural Network (QNN) following a data re-uploading approach Schuld et al. (2021); Casas and Cervera-Lierta (2023); Pérez-Salinas et al. (2020). As was proven in Schuld et al. (2021); Gil Vidal and Theis (2020), following an exponential data encoding approach leads the quantum model to represent a truncated Fourier series. Subsequently, each trigonometric term of the Fourier series undergoes quantum integration using Iterative Quantum Amplitude Estimation (IQAE) Grinko et al. (2021), an efficient version of Quantum Amplitude Estimation(QAE) Brassard et al. (2002).
@@
The Fourier decomposition enables the target function to be encodable with a minimum number of quantum arithmetic operations. At the same time, it takes advantage of the fact that the sine function is more suitable for integration in a quantum approach. The QNN is the central ingredient of QFIAE that offers a viable strategy to retain the potential quadratic speedup with respect to other quantum integration algorithms recently proposed, such as Fourier Quantum Monte Carlo Integration (FQMCI) Herbert (2022). The key idea of FQMCI is also to use Fourier series decomposition to approximate the integrand and then estimate each component separately using QAE. Nevertheless, it relies on certain assumptions regarding the acquisition of the Fourier coefficients, which may not hold in general. When these assumptions are not met, the quantum speedup might be wiped out. The QNN ensures a reliable extraction of the coefficients in a quantum way.
@@
The second crucial aspect of QFIAE involves leveraging the advantages of IQAE over QAE. QAE Brassard et al. (2002) is a quantum algorithm that estimates quantum state amplitudes using amplitude amplification, a generalization of Grover’s algorithm Grover (1997). This process enhances the probability of measuring a desired state over a nondesired state. However, QAE has inherent limitations, including its reliance on the resource-intensive Quantum Phase Estimation (QPE) subroutine Nielsen and Chuang (2011), which involves operations considered computationally expensive for current Noisy Intermediate Scale Quantum (NISQ) devices that may compromise the anticipated quadratic speedup promised by QAE. To address this challenge, IQAE replaces QPE with a classically efficient post-processing method, reducing the qubits and gates requirements while maintaining the asymptotic quadratic speedup.
@@
II II. QUANTUM INTEGRATION OF LOOP FEYNMAN INTEGRALS
@@
Among other advantages, this transformation provides a more intuitive understanding of the singular structure of loop integrals Buchta et al. (2014); Aguilera-Verdugo et al. (2019). In particular, the most remarkable property of LTD is the existence of a manifestly causal representation ”Aguilera-Verdugo et al. (2020); Aguilera-Verdugo et al. (2021); Ramírez-Uribe et al. (2021); Jesús Aguilera-Verdugo et al. (2021); Sborlini (2021); Torres Bobadilla (2021), i.e., an integrand representation where certain nonphysical singularities are absent and therefore yields integrands that are numerically more stable. Since the integration domain in LTD is Euclidean, not Minkowski, it offers additional advantages, both for analytic applications such as asymptotic expansions Driencourt-Mangin et al. (2018); Plenter and Rodrigo (2021), where the hierarchy of scales is well defined, and numerical applications Buchta et al. (2017); Driencourt-Mangin et al. (2022) because the number of loop integration variables is independent of the number of external particles. For example, at one loop the number of independent integration variables is always three, the number of spatial components of the loop momentum, although for certain kinematic configurations this number is reduced when the dependence on any of these variables is trivial. This is the case for tadpole and bubble diagrams at one loop. Moreover, LTD offers a unified framework for cross-section calculations, since the dual representation of loop integrals in Euclidean domains allows a direct combination of virtual and real contributions at the integrand level, resulting in a fully local cancellation of IR Hernandez-Pinto et al. (2016); Sborlini et al. (2016a, b); Prisco and Tramontano (2021); de Jesús Aguilera-Verdugo et al. (2021) and UV Driencourt-Mangin et al. (2018, 2019) singularities without the need for modifying the dimensions of space-time, such as in Dimensional Regularization (DREG) Bollini and Giambiagi (1972); ’t Hooft and Veltman (1972).
@@
II.1 A. Tadpole loop integral
The first Feynman loop integral we address is the one-loop tadpole integral
@@
is well defined in the four physical dimensions, where the integration measure reads $\int_{\boldsymbol{\ell}}=\int d^{3}\boldsymbol{\ell}/(2\pi)^{3}$. In Eq. (2), we have defined the on-shell energies $q_{1,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m^{2}-\imath 0}$ and $q_{{\rm UV},0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+\mu_{\rm UV}^{2}-\imath 0}$, where $\mu_{\rm UV}$ is the renormalization scale. The tadpole integral is one-dimensional because the integrand is independent of the solid angle, and only depends on the modulus of the three-momentum. The change of variable
$|\boldsymbol{\ell}|=\frac{m\,z}{1-z}\leavevmode\nobreak\ ,\qquad z=[0,1)\leavevmode\nobreak\ ,$ (3)
remaps the integration variable into a finite range where the Fourier series of the target function is defined.
To implement QFIAE in a real quantum computer we must consider that quantum computers operating in the NISQ era encounter diverse sources of noise, ranging from quantum effects like decoherence to hardware-specific errors, including gate, readout, and calibration errors. At the same time, working on the same hardware technology (superconducting qubits), we advocate for a hardware-agnostic implementation strategy. Specifically, we propose an approach where the two modules of QFIAE are implemented on two different quantum computers supplied by different providers. This enables a minimization of the impact of hardware-specific noise on the overall algorithmic performance.
@@
On the other hand, a 5-qubit IQAE algorithm has been executed using Qiskit Qiskit contributors (2023) on the 27-qubit IBMQ superconducting device ibmq_mumbai. To mitigate quantum noise during execution, we employed a pulse-efficient transpilation technique Earnest et al. (2021). Effectively reducing the number of two-qubit gates by harnessing the hardware-native cross-resonance interaction. While it has previously demonstrated promise in the context of Variational Quantum Circuits (VQC) Melo et al. (2023), it has not yet been extensively explored for fault-tolerant applications. Considering the positive results presented in Table 1, this work serves as an intriguing starting point for the applicability of this mitigation technique to fault-tolerant algorithms. Furthermore, we also applied two more error mitigation techniques, Dynamical Decoupling (DD) and Zero Noise Extrapolation (ZNE), using the Qiskit Runtime Estimator primitive IBMQ (2023).
Fourier IQAE ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$ $m=5,\mu_{\rm UV}=m/2$ ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$ $m=5,\mu_{\rm UV}=2m$
C C $-0.106$ $-0.258$
@@
Table 1: Renormalized tadpole integral ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$ on Qibo and IBMQ, as a function of the ratio of the renormalization scale $\mu_{\rm UV}$ to the mass $m$, which is fixed to $m=5$ GeV. The Fourier decomposition and IQAE integration are performed either with a quantum simulator (S) or on a real quantum device (Q). In the first raw both components are performed using classical standard methods (C).
In particular, the hardware implementation has been done by fixing the mass to $m=5$ GeV, and considering two values of the renormalization scale, $\mu_{\rm UV}=2m$ and $\mu_{\rm UV}=m/2$. The uncertainties on the integrals presented in Table 1 have been computed as the quadratic sum of the individual uncertainties provided by IQAE for each trigonometric component. These uncertainties are statistical in nature and are to be combined with the statistical uncertainties from the noise of the quantum devices employed in both the QNN training and IQAE integration modules of the QFIAE algorithm. Furthermore, there are also systematic uncertainties from the classical or quantum Fourier series truncation of the integrand function ${\cal A}^{(1,{\rm R})}_{1}(m;\mu_{\rm UV})$. Taking into account all statistical and systematic uncertainties the obtained results are in agreement with the analytical values within uncertainties.
The results presented in Table 1 show a relatively small deviation from the analytical value for both $\mu_{\rm UV}=m/2$ and $\mu_{\rm UV}=2m$. In particular, the agreement with the analytical values when both components of the algorithm are executed on a quantum computer (second to last row of Table 1) is better than $1.7$ standard deviations in all cases. This result represents a noteworthy achievement for the current state of the art in quantum computing technology. Moreover, considering that it is the first application of an end-to-end quantum algorithm executed on a quantum computer for estimating Feynman loop integrals. Furthermore, part of the deviation comes from the Fourier series approximation itself, as shown in the first row of Table 1, which constitutes proof of the robustness of the quantum algorithm presented.
II.2 B. Bubble loop integral
QFIAE is also effective in dealing with Feynman loop integrals with threshold singularities. Therefore, we consider the one-loop bubble integral:
@@
${\cal A}^{(1,{\rm R})}_{2}(p,m_{1},m_{2})=\int_{\boldsymbol{\ell}}\left[\frac{1}{x_{2}}\left(\frac{1}{\lambda^{+}}+\frac{1}{\lambda^{-}}\right)-\frac{1}{4(q_{{\rm UV},0}^{(+)})^{3}}\right]\leavevmode\nobreak\ ,$ (5)
where $x_{2}=\prod_{i=1,2}2q_{i,0}^{(+)}$, $\lambda^{\pm}=\sum_{i=1,2}q_{i,0}^{(+)}\pm p_{0}$, and the on-shell energies are given by $q_{i,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m_{i}^{2}-\imath 0}$ with $i\in\{1,2\}$, assuming the external momentum has vanishing spatial components, $p=(p_{0},{\bf 0})$. If $p_{0}^{2}<(m_{1}+m_{2})^{2}$ the integral is purely real. Otherwise, it gets an imaginary contribution from the unitary threshold singularity at $\lambda^{-}\to 0$, assuming $p_{0}>0$. To deal with this threshold singularity, it is convenient to introduce a contour deformation in the complex plane, in order to smooth the behavior of the function in the vicinity of the threshold without altering the result of the integral. We consider $m_{1}=m_{2}=m$, and proceed to estimate the integral using two different QNNs to fit separately the real and imaginary parts of the integrand and then integrate each Fourier series using IQAE. We have implemented QFIAE on two different quantum simulators. For the QNNs, we employed Pennylane Bergholm et al. (2018) whereas for IQAE, we utilized Qibo.
@@
II.3 C. Triangle loop integral
We now consider the one-loop three-point function, which corresponds to a triangle loop topology. The LTD representation of this integral is given by Aguilera-Verdugo et al. (2021)
@@
with $x_{3}=\prod_{i=1}^{3}2q_{i,0}^{(+)}$, where now the on-shell energies are $q_{1,0}^{(+)}=\sqrt{(\boldsymbol{\ell}+{\bf p}_{1})^{2}+m_{1}^{2}-\imath 0}$, and $q_{i,0}^{(+)}=\sqrt{\boldsymbol{\ell}^{2}+m_{i}^{2}-\imath 0}$ for $i\in\{2,3\}$. We work in the center of mass frame where ${\bf p}_{12}={\bf p}_{1}+{\bf p}_{2}=0$, $p_{12,0}=p_{1,0}+$ $p_{2,0}=\sqrt{s}$ and the external momenta $p_{1}$ and $p_{2}$ are back-to-back along the $z$ axis. The causal denominators are
@@
In this work, we have for the first time successfully implemented a quantum algorithm for evaluating Feynman loop integrals in quantum hardware. Our approach is based on the nonconventional LTD representation of Feynman loop integrals and the QFIAE quantum algorithm which provides a theoretical quadratic quantum speedup in the number of queries of the probability distribution function.
@@
We have provided a clear indication of how the optimization of a quantum circuit, at the level of pulses and the application of combined error mitigation techniques, enable the collection of accurate results. Specifically, we employed QFIAE to numerically integrate the one-dimensional tadpole Feynman loop integral entirely on real quantum hardware provided by Qibo and IBM. Moreover, loop Feynman integrals with more external legs and dimensions have also been integrated using QFIAE on quantum simulators across various kinematical regions, circumventing different threshold singularities using a contour deformation. The obtained results demonstrate the remarkable performance of QFIAE in estimating these integrals.
@@
Overall, our study highlights the potential of quantum Monte Carlo integrators, particularly the QFIAE quantum algorithm, in integrating Feynman loop integrals. These findings, some of them obtained from quantum superconducting devices, serve as an initial step toward the search for a quantum method capable of speeding up the computationally expensive task of evaluating Feynman multiloop-multileg integrals. Moreover, they offer a proof of concept for the method’s applicability to wider challenges across various domains that require Monte Carlo integration, such as finance, artificial intelligence and other physical sciences.
When considering the scalability of QFIAE to higher dimensions, it is important to analyze its two key components separately. Regarding the IQAE integration, the use of Fourier decomposition allows us to express a multidimensional integral as a product of individual integrals containing Fourier terms with independent integration variables. Consequently, regardless of the dimensionality, IQAE is always applied to one-dimensional trigonometric integrands after Fourier decomposition. This ensures its efficient extension to multidimensional functions. In relation to the QNN fitting the target function, universality has been theoretically demonstrated in Schuld et al. (2021) for exponential encoding and a data re-uploading scheme. This constitutes a theoretical proof that the first component of QFIAE is also scalable to higher dimensions.
@@
Still, the main challenge lies in identifying a suitable Ansatz capable of ensuring the trainability of multidimensional functions, which is essential for addressing higher-loop integrals. While universality has proven effective in lower dimensions, extending it to higher dimensions remains an open problem. In this work, we achieved a successful fit for one and two-dimensional functions. However, the search for a viable Ansatz in higher dimensions which involves a tradeoff between trainability and expressivity of the variational quantum circuit is left for future work.
Progressing in the previous limitation would assure, according to the algorithm proposed in this article, a significant noteworthy advancement over classical Monte Carlo methods for computing these integrals, presenting new opportunities for unlocking previously inaccessible levels of precision in theoretical predictions for high-energy physics at colliders.
IV Acknowledgments